# Charactarisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v1

@@ 3. Characterisation of distal actions on the space of one-parameter subgroups
Let ${\mathcal{K}}$ be the connected component of the identity in the closed subgroup generated by $\overline{T}$ in ${\rm Aut}(G/M)$, then $(\overline{T})^{k}\in{\mathcal{K}}$ for some $k\in{\mathbb{N}}$. Now by Lemma 3.9, $\overline{T}$ keeps a maximal torus (say), $A$ of $G/M$ invariant. Let $K=\pi^{-1}(A)$. Then $K$ is a maximal torus of $G$ and $T^{k}(K)=K$.
Note that $K$ is compact, connected and abelian. Moreover, as $T$ acts distally on $G$, so does $T^{k}$ and the restriction of $T^{k}$ to $K$ is also distal. Arguing as above for $K$ instead of $M$, we get that for some $m\in{\mathbb{N}}$, $T^{m}|_{K}$ is unipotent, and as it acts distally on ${\rm Sub}^{p}_{K}$, it acts trivially on $K$ (by Theorem 1.3). Replacing $T$ by $T^{m}$, we may assume that $T$ fixes every element of $K$, i.e. $T\in F_{K}(G)=\{\psi\in{\rm Aut}(G)\mid\psi(x)=x\mbox{ for all }x\in K\}$. We know that $F_{K}(G)$ is almost algebraic (cf. [15]). As $T$ is distal and $T\in F_{K}(G)$, for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathcal{B}}\times{\mathcal{U}}$, where ${\mathcal{B}}$ is a compact group and ${\mathcal{U}}$ consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now we have that $T=\phi\circ S=S\circ\phi$ such that $\phi\in{\mathcal{B}}$ and $S\in{\mathcal{U}}$. As ${\mathcal{B}}$ is compact, we have that $S$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.3, $S={\rm Id}$. Thus $T=\phi\in{\mathcal{B}}$ and (5) holds. ∎
@@
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Suppose ${\mathcal{H}}$ acts distally on the largest compact connected central subgroup of $G$. Then so does $\overline{{\mathcal{H}}}$, which is a closed subgroup of ${\rm Aut}(G)$. Note that ${\mathcal{H}}$ acts distally on a compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$ (both the above statements follow from [16]). Also (7) holds for ${\mathcal{H}}$ if and only if it holds when ${\mathcal{H}}$ is replaced by $\overline{{\mathcal{H}}}$, as ${\mathcal{K}}{\mathcal{D}}$ is closed in ${\rm GL}(d,{\mathbb{R}})$ for ${\mathcal{K}},{\mathcal{D}}$ as in (7). Therefore to prove the assertions, we can assume that ${\mathcal{H}}$ is closed. Note that every compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [39], Lemma 2.2). Now $(6)\implies(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious.
@@
First suppose that $G$ is not a vector group. We show that $(1)\implies(6)$. As in the hypothesis, Suppose (1) holds, i.e. every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.4, every $T\in{\mathcal{H}}$ is contained in a compact subgroup of ${\rm Aut}(G)$. Recall that ${\rm Aut}(G)$ is identified with a closed subgroup of ${\rm GL}({\mathfrak{g}})$ via the map $T\to{\rm d\,}T$, where ${\mathfrak{g}}$ is the Lie algebra of $G$ and it is a $d$-dimensional vector space for some $d\in{\mathbb{N}}$ (cf. [9], see also [21]). Therefore, ${\mathcal{H}}$ is also a closed subgroup of ${\rm GL}({\mathfrak{g}})$ and the latter is isomorphic to ${\rm GL}(d,{\mathbb{R}})$, and we get by Proposition 2 of [38] that ${\mathcal{H}}$ is compact. Thus (6) holds and $(1-6)$ are equivalent. Therefore, statements (a), (b) and (c) hold in case $G$ is not a vector group.
@@
Now suppose $G={\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$. Suppose (3) holds, i.e. ${\mathcal{H}}$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$. Then every $T\in\rm{(NC)}$. Since $G$ belongs to the class ${\mathcal{C}^{\prime}}$, by Theorem 4.1 of [40], $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. Since ${\mathcal{H}}$ is closed, arguing as above using Proposition 2 of [38], we get that ${\mathcal{H}}$ is compact. Hence (6) holds. Thus $(3-6)$ are equivalent and hence (c) holds when $G$ is a vector group. Now we show that (1), (2) and (7) are equivalent. Suppose (7) holds. Note that ${\mathcal{D}}=\{r\,{\rm Id}\in{\rm GL}(d,{\mathbb{R}})\mid r\in{\mathbb{R}}\setminus\{0\}\}$. Therefore, ${\mathcal{D}}$ acts trivially on ${\rm Sub}^{p}_{G}$ and the action of ${\mathcal{K}}{\mathcal{D}}$ on ${\rm Sub}^{p}_{G}$ is same as that of the compact group ${\mathcal{K}}$ on ${\rm Sub}^{p}_{G}$. Hence the action of ${\mathcal{K}}{\mathcal{D}}$, and hence, the action of ${\mathcal{H}}$ on ${\rm Sub}^{p}_{G}$ is distal. Thus $(7)\implies(2)$. We have already noted that $(2)\implies(1)$.
@@
Corollary 3.10 is a direct consequence of Theorem 1.5 in case $G$ is not a vector group. If $G$ is a vector group, then as ${\mathcal{H}}$ acts distally on $G$, the eigenvalues of every element in ${\mathcal{H}}$ have absolute value 1 (cf. [1]), and this together with the statement (7) in Theorem 1.5 is equivalent to the statement that $\overline{{\mathcal{H}}}$ is compact. We omit the details.
@@
Suppose (1) holds; i.e. every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$. If $G$ is either compact or a vector group, then (5) holds trivially. Now suppose $G$ neither compact nor a vector group. Since every inner automorphism acts trivially on the center of $G$, we have by Theorem 1.2 that it acts distally on $G$. Therefore, $G$ is distal (cf. [35]), i.e. the action of ${\rm Inn}(G)$ on $G$ is distal. It follow from Theorem 9 of [35] and Corollary 2.1 (ii) of [26] that $G/R$ is compact, where $R$ is the radical of $G$. Let $N$ be the nilradical of $G$ and let $x\in N$. Then ${\rm inn}(x)$ is unipotent, and by Theorem 1.3, ${\rm inn}(x)={\rm Id}$. This implies that $x$ is central in $G$ for every $x\in N$. Then $N$ is abelian and central in $G$. Now for the radical $R$, we know that $[R,R]\subset N$ is central in $G$, and hence $R$ is nilpotent and it is same as $N$ and it is abelian and central in $G$. Now $G/N$ is compact, where $N$ is central in $G$, i.e. $N={\mathbb{R}}^{n}\times M$, for some $n\in{\mathbb{N}}$, where $M$ is compact. Here, $V={\mathbb{R}}^{n}$ is a vector subgroup which is central. By Lemma 3.7 of [24], $G=K\times{\mathbb{R}}^{n}$, where $K$ is the maximal compact subgroup of $G$. Thus (5) holds. ∎
The distal actions of automorphisms $T$ of a connected Lie group $G$ on ${\rm Sub}_{G}$ and its closed $T$-invariant subspaces were studied if $G$ is in class ${\mathcal{C}^{\prime}}$ or $T$ is unipotent; here we have studied a more general case when $T$ acts distally on the maximal central torus of $G$. We have also studied the action of subgroups of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$ for a connected Lie group $G$ under a condition that the action on the maximal central torus is distal. It would be interesting to characterise general automorphisms of $G$ which act distally on ${\rm Sub}_{G}$ or its closed invariant subspace for any connected Lie group $G$ without the condition involving the action on the central torus.

### v2

@@ 4. Proofs of the main results
Let ${\mathcal{K}}$ be the connected component of the identity in the closed subgroup generated by $\overline{T}$ in ${\rm Aut}(G/M)$, then $(\overline{T})^{k}\in{\mathcal{K}}$ for some $k\in{\mathbb{N}}$. Now by Lemma 4.3, $(\overline{T})^{k}$ keeps a maximal torus (say) $A$ of $G/M$ invariant and it acts trivially on $A$. Let $K=\pi^{-1}(A)$. Then $K$ is a maximal torus of $G$ and $T^{k}(K)=K$.
Note that $K$ is compact, connected and abelian. Moreover, as $T$ acts distally on $G$, so does $T^{k}$ and the restriction of $T^{k}$ to $K$ is also distal. Arguing as above for $K$ instead of $M$, we get that for some $m\in{\mathbb{N}}$, $T^{m}|_{K}$ is unipotent and, as it acts distally on ${\rm Sub}^{p}_{K}$, it acts trivially on $K$ (by Theorem 1.3). Without loss of any generality, we may replace $T$ by $T^{m}$ and assume that $T$ fixes every element of $K$, i.e. $T\in F_{K}(G):=\{\varrho\in{\rm Aut}(G)\mid\varrho(x)=x\mbox{ for all }x\in K\}$. We know that $F_{K}(G)$ is almost algebraic (cf. [15]). As $T$ acts distally on $G$ and $T\in F_{K}(G)$, for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathcal{B}}\times{\mathcal{U}}$, where ${\mathcal{B}}$ is a compact group and ${\mathcal{U}}$ consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now we have that $T=\phi\circ S=S\circ\phi$ such that $\phi\in{\mathcal{B}}$ and $S\in{\mathcal{U}}$. As ${\mathcal{B}}$ is compact, we have that $S$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.3, $S={\rm Id}$. Thus $T=\phi\in{\mathcal{B}}$ and (5) holds. ∎
@@
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Suppose ${\mathcal{H}}$ acts distally on the largest compact connected central subgroup of $G$. Then so does $\overline{{\mathcal{H}}}$. Note that ${\mathcal{H}}$ acts distally on a compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$ (both the above statements follow from Theorem 1 in [16]). Also (7) holds for ${\mathcal{H}}$ if and only if it holds when ${\mathcal{H}}$ is replaced by $\overline{{\mathcal{H}}}$, as ${\mathcal{K}}{\mathcal{D}}$ is closed in ${\rm GL}(d,{\mathbb{R}})$ for ${\mathcal{K}},{\mathcal{D}}$ as in (7). Therefore to prove the assertions, we may assume that ${\mathcal{H}}$ is closed. Note that every compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [41], Lemma 2.2). Now $(6)\implies(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious.
@@
First suppose that $G$ is not a vector group. We show that $(1)\implies(6)$. Suppose (1) holds, i.e. every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.4, every $T\in{\mathcal{H}}$ is contained in a compact subgroup of ${\rm Aut}(G)$. Recall that ${\rm Aut}(G)$ is identified with a closed subgroup of ${\rm GL}({\mathfrak{g}})$ via the map $T\to{\rm d\,}T$, where ${\mathfrak{g}}$ is the Lie algebra of $G$ and it is a $d$-dimensional vector space for some $d\in{\mathbb{N}}$ (cf. [9], see also [22]). Therefore, ${\mathcal{H}}$ is also a closed subgroup of ${\rm GL}({\mathfrak{g}})$ and the latter is isomorphic to ${\rm GL}(d,{\mathbb{R}})$, and we get by Theorem 1.1 of [17] (see also Proposition 2 of [39]) that ${\mathcal{H}}$ is compact. Thus (6) holds and $(1-6)$ are equivalent. Thus, if $G$ is not a vector group, the assertions in the theorem hold for this case.
@@
Now suppose $G={\mathbb{R}}^{d}$ for some $d\in{\mathbb{N}}$. Suppose (3) holds, i.e. ${\mathcal{H}}$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$. Then $T\in\rm{(NC)}$ for every $T\in{\mathcal{H}}$. Since $G$ belongs to the class ${\mathcal{C}^{\prime}}$, by Theorem 4.1 of [42], $T$ is contained in a compact subgroup of ${\rm Aut}(G)$ for every $T\in{\mathcal{H}}$. Since ${\mathcal{H}}$ is closed, arguing as above using Theorem 1.1 of [17] or Proposition 2 of [39], we get that ${\mathcal{H}}$ is compact. Hence (6) holds. Thus $(3-6)$ are equivalent when $G$ is a vector group. Now we show that (1), (2) and (7) are equivalent. Suppose (7) holds. Note that ${\mathcal{D}}=\{r\,{\rm Id}\in{\rm GL}(d,{\mathbb{R}})\mid r\in{\mathbb{R}}\setminus\{0\}\}$. Therefore, ${\mathcal{D}}$ acts trivially on ${\rm Sub}^{p}_{G}$ and the action of ${\mathcal{K}}{\mathcal{D}}$ on ${\rm Sub}^{p}_{G}$ is same as that of the compact group ${\mathcal{K}}$ on ${\rm Sub}^{p}_{G}$. Therefore, the action of ${\mathcal{K}}{\mathcal{D}}$, and hence, the action of ${\mathcal{H}}$ on ${\rm Sub}^{p}_{G}$ is distal. Thus $(7)\implies(2)$. We have already noted that $(2)\implies(1)$.
Now we show that $(1)\implies(7)$. Suppose every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. Let $T\in{\mathcal{H}}$ be fixed. By Theorem 1.1 (1), $T\in{\mathcal{B}}{\mathcal{D}}$, for some compact group ${\mathcal{B}}\subset{\rm GL}(d,{\mathbb{R}})$. Let ${\mathcal{H}}_{1}:=\overline{[{\mathcal{H}},{\mathcal{H}}]}$. Then ${\mathcal{H}}_{1}\subset{\mathcal{H}}$ as the latter is closed. Now let $T\in{\mathcal{H}}_{1}$. Then $\det T=1$. From above we have that $T=\phi\,\circ\,r\,{\rm Id}$ for some $\phi\in{\mathcal{B}}$ and some $r\in{\mathbb{R}}\setminus\{0\}$. Since ${\mathcal{B}}$ is a compact group, $\det\phi=\pm 1$, and hence $r=\pm 1$. Therefore, $T$ generates a relatively compact group, for every $T\in{\mathcal{H}}_{1}$. As ${\mathcal{H}}_{1}$ is a closed subgroup of ${\rm GL}(d,{\mathbb{R}})$ by Theorem 1.1 of [17] or Proposition 2 of [39], ${\mathcal{H}}_{1}$ is compact. Now ${\mathcal{H}}_{1}$ is a compact normal subgroup of ${\mathcal{H}}$. Let $N({\mathcal{H}}_{1})$ be the normaliser of ${\mathcal{H}}_{1}$ in ${\rm GL}(d,{\mathbb{R}})$. Then $N({\mathcal{H}}_{1})$ is a closed algebraic subgroup of ${\rm GL}(d,{\mathbb{R}})$ and ${\mathcal{H}},{\mathcal{D}}\subset N({\mathcal{H}}_{1})$. Moreover, $N({\mathcal{H}}_{1})\subset N({\mathcal{H}}_{1}{\mathcal{D}})$. Being algebraic, $N({\mathcal{H}}_{1})$ has finitely many connected components, and hence so does $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$. In particular any closed abelian subgroup of $N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ is compactly generated (cf. [24]). Let $\rho:N({\mathcal{H}}_{1})\to N({\mathcal{H}}_{1})/{\mathcal{H}}_{1}{\mathcal{D}}$ be the natural projection. Then $\rho({\mathcal{H}})$ is abelian, and hence so is $\overline{\rho({\mathcal{H}})}$. Therefore, $\overline{\rho({\mathcal{H}})}$ is compactly generated, and it has a unique maximal compact subgroup, say ${\mathcal{C}}$.
@@
Corollary 4.4 is a direct consequence of Theorem 1.5 in case $G$ is not a vector group. If $G$ is a vector group, then as ${\mathcal{H}}$ acts distally on $G$, the eigenvalues of every element in ${\mathcal{H}}$ have absolute value 1 (cf. [1]), and this, together with the statement (7) in Theorem 1.5, is equivalent to the statement that $\overline{{\mathcal{H}}}$ is compact. We omit the details.
@@
Suppose (1) holds; i.e. every inner automorphism of $G$ acts distally on ${\rm Sub}^{p}_{G}$. If $G$ is either compact or a vector group, then (5) holds trivially. Now suppose $G$ neither compact nor a vector group. Since every inner automorphism acts trivially on the center of $G$, we have by Theorem 1.2 that it acts distally on $G$. Therefore, $G$ is distal (cf. [36]), i.e. the action of ${\rm Inn}(G)$ on $G$ is distal. It follows from Theorem 9 of [36] and Corollary 2.1 (ii) of [27] that $G/R$ is compact, where $R$ is the radical of $G$. Let $N$ be the nilradical of $G$ and let $x\in N$. Then ${\rm inn}(x)$ is unipotent, and by Theorem 1.3, ${\rm inn}(x)={\rm Id}$. This implies that $x$ is central in $G$ for every $x\in N$. Then $N$ is abelian and central in $G$. Now for the radical $R$, we know that $[R,R]\subset N$ is central in $G$, and hence $R$ is nilpotent and it is same as $N$ and it is abelian and central in $G$. Now $G/N$ is compact, where $N$ is central in $G$, i.e. $N={\mathbb{R}}^{n}\times M$, for some $n\in{\mathbb{N}}$, where $M$ is compact. Here, $V:={\mathbb{R}}^{n}$ is a vector subgroup which is central. By Lemma 3.7 of [25], $G=K\times{\mathbb{R}}^{n}$, where $K$ is the maximal compact subgroup of $G$. Thus (5) holds. ∎
The distal actions of automorphisms $T$ of a connected Lie group $G$ on ${\rm Sub}^{a}_{G}$ were studied earlier if $G$ is in class ${\mathcal{C}^{\prime}}$ or $T$ is unipotent. Here we have studied a more general case when $T$ acts distally on the maximal central torus of $G$ and the action is considered on ${\rm Sub}^{p}_{G}$, a smaller subspace of ${\rm Sub}^{a}_{G}$. More generally, we have also studied the action of subgroups ${\mathcal{H}}$ of ${\rm Aut}(G)$ on ${\rm Sub}^{p}_{G}$ under a condition that the ${\mathcal{H}}$-action on the maximal central torus of $G$ is distal. This latter condition is satisfied if ${\mathcal{H}}$ is contained in ${\rm Aut}(G)^{0}$ or more generally, if it is contained in an almost algebraic subgroup of ${\rm Aut}(G)$.
The following example of a connected solvable Lie group $G$, known as the Walnut group, is not covered by Theorem 1.1 or any earlier theorems for the distal action of a general automorphism on ${\rm Sub}^{a}_{G}$, as it has nontrivial central torus and ${\rm Aut}(G)$ is not almost algebraic (see [15]). But it is covered by theorems proven here.
Example 4.6.
Let $\mathbb{H}$ be the 3-dimensional Heisenberg group of $3\times 3$ strictly upper triangular matrices. There is a canonical action of $SL(2,{\mathbb{R}})$ on $\mathbb{H}$ which fixes elements of its center, which is isomorphic to ${\mathbb{R}}$. Let $N=\mathbb{H}/D$, where $D$ is a discrete central subgroup isomorphic to ${\mathbb{Z}}$. Then $N$ is a 2-step nilpotent group and the ${\rm SL}(2,{\mathbb{R}})$-action on $\mathbb{H}$ carries over to the action on $N$. Let $G=SO(2,{\mathbb{R}})\ltimes N$, the Walnut group. Since the center of $G$ is isomorphic to ${\mathbb{S}}^{1}$, ${\rm Aut}(G)$ acts distally on it. Thus Theorems 1.2 and 1.4 hold for the action of any $T\in{\rm Aut}(G)$ and Theorem 1.5 holds for any subgroup $H$ of ${\rm Aut}(G)$.
One can also take a non-solvable group $G=SL(2,{\mathbb{R}})\ltimes N$, for $N$ mentioned as above, whose automorphism group is almost algebraic (see [15]). Theorems mentioned in the above example hold also for this group. This group $G$ is mentioned in Remark 4.8 of [42], although it has a maximal torus of dimension 2, ${\rm Aut}(G)$ is almost algebraic (see [15]), and the conclusion in Theorem 4.1 of [42] mentioned in the remark holds for this $G$.
It would be interesting to characterise general automorphisms of any connected Lie group $G$ which act distally on ${\rm Sub}_{G}$ or its closed invariant subspace without any condition involving the action on the central torus.